OpenStudy (anonymous):

Hi good day I'm Joshua James A. Bravante. I'm taking up BS Mathematics minor in Statistics, from Southern Luzon State University Lucban, Quezon Philippines. I need some help for my project in Mathematical Analysis III that we need to exchange idea in particular topic. I choose vector in three dimensional . Can you help me?

OpenStudy (anonymous):

can you be specific ? Where do you need help in 3D ?

OpenStudy (anonymous):

specifically in rectangular coordinate system

OpenStudy (shubhamsrg):

Are you asking for my help ?

OpenStudy (anonymous):

yes, please,

OpenStudy (shubhamsrg):

wow! Thats the first time in my life! :')

OpenStudy (shubhamsrg):

I'll see what I can do !

OpenStudy (anonymous):

:'( so i am leaving

OpenStudy (anonymous):

We need to exchange ideas about a particular topic, 3D, particularly rectangular coordinate system.

OpenStudy (anonymous):

i think i can help u

OpenStudy (anonymous):

Let's talk about 3 dimensional, particularly in rectangular coordinate system. What can you say about that?

OpenStudy (anonymous):

based on the definition of three dimensional the set of real numbers is called the three dimensional number space , denoted R cube.Each ordered triple pairs (x,y,z) is called a point in three dimensional number space.

terenzreignz (terenzreignz):

Right...

terenzreignz (terenzreignz):

And each vector in 3d corresponds to a point in the Cartesian space. (R cubed)

OpenStudy (anonymous):

Yes then to represent in a geometric three dimensional , we need to consider that directed distance of a point from three mutually perpendicular plane. How about if recrangular coordinates system.?

terenzreignz (terenzreignz):

I'm not sure what that question meant, but something about the 3d vectors, they're the ones with the cross-product defined on them.

OpenStudy (anonymous):

what can you share about cross product in 3D

terenzreignz (terenzreignz):

Cross product of two vectors in 3d generates a vector which is perpendicular (usually say orthogonal) to both vectors that were cross-product---ed XD Also, cross product is not commutative, unlike dot product. Two vectors in 3d are parallel if and only if their cross product is the zero vector.

OpenStudy (anonymous):

Yes based on the theorem If A and B are two vectors in V sub 3, then the vector A, X, B is orthogonal to both A and B.

terenzreignz (terenzreignz):

I've run short of what I know about vectors in 3d :)

OpenStudy (anonymous):

okay, now lets move on to the next interesting topic... quadric surface, what do you know about that? :)

terenzreignz (terenzreignz):

Quadric surfaces are the extension of conic sections in 2d. They are the Ellipsoid, the Elliptic Paraboloid, Elliptic Hyperboloid (of one or two sheets) and the Hyperbolic Paraboloid

OpenStudy (anonymous):

how to determine the characteristics of quadric surface coresponding to a general equation?

terenzreignz (terenzreignz):

I haven't memorised them, sorry :) try this http://en.wikipedia.org/wiki/Quadric

OpenStudy (anonymous):

okay i will just refer to the link you had sent me... i have this topic which i want to know more about it deeply. it is all about parametric equation.

terenzreignz (terenzreignz):

Well, it's when both x and y are functions of some parameter t. x = f(t) y = g(t) For instance x = 3t + 5 y = 8 - 3t These are parametric equations for the line y =13 - x This can easily be seen, as you just solve for t in terms of x, and substitute, however, something more interesting...

terenzreignz (terenzreignz):

x = cos t y = sin t It might not be intuitively obvious, but actually, these are the parametric equations for the unit circle!

terenzreignz (terenzreignz):

Because 1 = cos²t + sin²t = (cos t)² + (sin t)² = x² + y ² 1 = x² + y²

OpenStudy (anonymous):

can you decipher that for me.

terenzreignz (terenzreignz):

You know that the equation of a unit circle is x² + y ² = 1 Now, parametric equations were x = cos t y = sin t From the Pythagorean identities, it follows that cos²t + sin²t = 1 Substituting x for cos t and y for sin t, we get x² + y² = 1 Which is precisely the equation of a unit circle.

OpenStudy (anonymous):

can you please help me?